OpenStudy (anonymous):

convert .98 to percent

OpenStudy (anonymous):

98 %

OpenStudy (anonymous):

.98 *100

OpenStudy (anonymous):

0.98 = 98/100 = 98%

OpenStudy (anonymous):

Just multiply bu 100

OpenStudy (anonymous):

ok what about 1 1/4 to a percent

OpenStudy (anonymous):

\[.98 \]\[= \text{98 hundredths} \]\[= \frac{98}{100} \]\[= \text{ 98 per 100} \]\[= \text{98 per centum} \]\[= \text{98 percent} \]\[=98 \%\]

OpenStudy (anonymous):

\[\text{1 and 1/4}\]\[=1 + 1/4\]\[=1 + .25\]\[=1.25\]\[=\frac{125}{100}\]\[=\text{125 per 100}\]\[=\text{125 per centum}\]\[=\text{125 percent}\]\[=125\%\]